# On the Erdős distance problem

### v8

@@ 3. Application to the Erdős unit distance and distinct distance conjecture
$=C\frac{\sqrt{k}}{2}n^{1+o(1)}$
for some $C>0$, by an application of Lemma 2.7 which establishes the claimed lower bound for the construction. ∎
It is important to point out that the lower estimate for the construction provided in Theorem 3.1 was achieved by counting not all possible unit distances but only unit distance that correspond to compression gaps of unit length. We state the second theorem as an application, which gives lower bound for the number of distinct distances that can be formed from $n$ points in the Euclidean space $\mathbb{R}^{k}$ for all $k\geq 2$.
@@
First, we set $m:=m(k)=o(1)$ as $k\longrightarrow\infty$ and carefully choose $n$ points $\vec{x}_{j}$ for $1\leq j\leq n$ in $\mathbb{R}^{k}$ such that $\lfloor\frac{n}{2}\rfloor$ of these points also has their image points under compression. That is, for each $\vec{x}_{j}$ we also include $\mathbb{V}_{m}[\vec{x}_{j}]$. Next for $\lfloor\frac{n}{2}\rfloor$ of those points we make the assignment $\mathrm{sup}(x_{j_{i}})=n^{1-\frac{2}{k}+\epsilon}$ for any small $\epsilon:=\epsilon(i)>0$ and $\mathrm{inf}(x_{j_{i}})\geq 1$. This ensures that
@@
for any small $\epsilon:=\epsilon(i)>0$. Now, we let $\{d_{j}:d_{j}=||\vec{x_{s}}-\vec{y_{t}}||,~{}d_{j}\neq d_{i},~{}1\leq s,t\leq n,~{}\vec{x},\vec{y}\in\mathbb{R}^{k}\}=\mathcal{R}$, then we notice that
$\#\mathcal{R}$$\geq\#\bigg{\{}d_{j}:d_{j}=||\vec{x_{s}}-\vec{y_{t}}||,~{}d_{j}\neq d_{i},i\neq j,~{}1\leq s,t\leq n,~{}\vec{x},\vec{y}\in\mathbb{R}^{k},~{}\mathrm{sup}(d_{j})=n^{1-\frac{2}{k}+o(1)}\bigg{\}}$
#$\geq\#\bigg{\{}d_{j}:d_{j}=\mathcal{G}\circ\mathbb{V}_{m}[\vec{x_{j}}],~{}d_{j}\neq d_{i},~{}1\leq j\leq\frac{n}{2},~{}\mathrm{sup}(d_{j})=n^{1-\frac{2}{k}+o(1)},~{}\vec{x_{j}}\in\mathbb{R}^{k},x_{j_{s}}~{}(1\leq s\leq k)$
$\geq 1,~{}\mathbb{V}[\vec{x_{j}}]=\vec{x_{t}}\bigg{\}}$
@@
and the claimed lower bound follows by Lemma 2.7 for this construction. ∎
It needs to be said that the result in Theorem 3.2 can be viewed as providing an alternate solution to the Erdős distinct distance problem, that takes into consideration the dimension of the space in which the points reside. The lower bound of this type, it has to be said, exists in the literature (See [1]). But the method employed is completely different from the one we have used here. Theorem 3.1 and Theorem 3.2 can be considered as a generalization of the solution to both versions of the Erdős distance problem to any euclidean space of dimension $k\geq 2$. In particular we have the following theorems as consequences of the main results of this paper.

### v10

@@ 3. Application to the Erdős unit distance and distinct distance conjecture
$=C\frac{\sqrt{k}}{2}n^{1+\frac{1}{(\log n)^{1+\epsilon}}}$
for some $C>0$, by application of the lemma 2.11. This establishes the claimed lower bound for the construction. ∎
It is important to note that the lower estimate for the construction provided in Theorem 3.1 was achieved by counting not all possible unit distances, but only the unit distance that corresponds to compression gaps of unit length. We state the second theorem as an application, which gives a lower bound for the number of distinct distances that can be formed from $n$ points in Euclidean space $\mathbb{R}^{k}$ for all $k\geq 2$.
@@
First, we set $m:=m(k)=o(1)$ as $k\longrightarrow\infty$ and carefully choose $n$ points $\vec{x}_{j}$ for $1\leq j\leq n$ in $\mathbb{R}^{k}$ such that $\lfloor\frac{n}{2}\rfloor$ of these points also have their image points under compression. That is, for each $\vec{x}_{j}$, we also include $\mathbb{V}_{m}[\vec{x}_{j}]$. Next, for $\lfloor\frac{n}{2}\rfloor$ of those points, we make the assignment $\mathrm{sup}(x_{j_{i}})=n^{1-\frac{2}{k}+\epsilon}$ for any small $\epsilon:=\epsilon(i)>0$ and $\mathrm{inf}(x_{j_{i}})\geq 1$. This ensures that
@@
for any small $\epsilon:=\epsilon(i)>0$. Now, we let
[…]
then
[…]
$\geq\#\bigg\{d_{j}:d_{j}=\mathcal{G}\circ\mathbb{V}_{m}[\vec{x_{j}}],~d_{j}\neq d_{i},~1\leq j\leq\frac{n}{2},~\mathrm{sup}(d_{j})=n^{1-\frac{2}{k}+o(1)},$
$\vec{x_{j}}\in\mathbb{R}^{k},x_{j_{s}}\geq 1,~(1\leq s\leq k),~\mathbb{V}[\vec{x_{j}}]=\vec{x_{t}}\bigg\}$
@@
where we have used the lemma 2.11, and the claimed lower bound follows for this construction. ∎
4. Conclusion
It needs to be said that the result in Theorem 3.2 can be viewed as providing an alternate solution to the Erdős distinct distance problem which takes into consideration the dimension of the space in which the points reside. The lower bound of this type exists in the literature (see, e.g, [1]). However, the method used is completely different from the one we have used here. Theorem 3.1 and Theorem 3.2 can be considered as a generalization of the solution to both versions of the Erdős distance problem to any euclidean space of dimension $k\geq 2$. In particular, we have the following theorems as consequences of the main results of this paper.